# IBNorm: Information-Bottleneck Inspired Normalization for Representation Learning

### v1

@@ E.4 Proof for Lemma 5
$|\mathbb{E}[p(\tilde{T}_{1}|X)]-\mathbb{E}[p(\tilde{T}_{1}^{\prime}|X)]|\leq\delta,\quad|\mathbb{E}[p(\tilde{T}_{1}|Y)]-\mathbb{E}[p(\tilde{T}_{1}^{\prime}|Y)]|\leq\delta,$
$|\mathbb{V}(p(\tilde{T}_{1}|X))-\mathbb{V}(p(\tilde{T}_{1}^{\prime}|X))|\leq\delta,\quad|\mathbb{V}(p(\tilde{T}_{1}|Y))-\mathbb{V}(p(\tilde{T}_{1}^{\prime}|Y))|\leq\delta.$
@@ E.5 Proof for Corollary 2
With probability at least $1-\delta$ over training set $S$ of size $M$, the generalization error of a network $f_{\text{o}}$ satisfies
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U_{l}=\sqrt{\tfrac{I(X;\tilde{T}_{l}^{\text{o}}|Y)+C(S,f_{\text{o}})+\log(\frac{2}{\delta})}{M}},$
where $\tilde{T}^{o}_{l}$ is the intermediate representation at layer $l$ before normalization representation recovery $\eta$ and $C(S,f_{\text{o}})$ is a term depending on the network $f_{\text{o}}$ and the training set $S$. Here $f_{\text{o}}$ can be the network $f_{\text{S}}$ using standard normalization like LN and BN, or our network $f_{\text{IB}}$ with our IBNorm.
@@
Theorem 2 in (Kawaguchi et al., 2023) states that, for any $\delta>0$, with probability at least $1-\delta$ over the training set $S$, the following generalization bound holds:
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U_{l}=\mathcal{O}\Big(\sqrt{\tfrac{I(X;\tilde{T}_{l}^{\text{o}}|Y)+C(S,f_{\text{o}})}{M}}\Big),$
where $\tilde{T}_{l}$ is the intermediate representation at layer $l$ before normalization representation recovery $\eta$ and $C(S,f_{\text{o}})$ is a term depending on the network $f_{\text{o}}$ and the training set $S$. Based on this theorem, we can derive a mutual information related bound as shown in Corollary 2. ∎
Remark. In our experiments, all models are trained on the same dataset, so any dataset-dependent contribution to the generalization bound can be treated as invariant across different normalization schemes. Although the output weight of $f_{\text{o}}$ during the training process can be thought as a random variable, their information content can be controlled in the training configuration, for instance by injecting noise drawn from a chosen distribution, with variance modulated from zero (allowing, in theory, full memorization of the training set) to a level large enough that no information remains (Achille & Soatto, 2018). Hence, we focus the role of normalization in our analysis: we examine how it shapes the information content of the intermediate representations and thereby affects generalization.
@@ F.1 LLM Experiments
For all LLaMA models, we follow the setup in (Zhao et al., 2024) and pretrain the vanilla LLaMA series models from scratch on the C4 dataset. We set the maximum input sequence length to 512 and output sequence length to 256, and adopt the bfloat16 precision. For the LLaMA-60M model, we use a learning rate of $0.001$, mini-batch size of 64 with gradient accumulation to a global batch size of 512, 20,000 training steps with 1,000 warm-up steps, and the AdamW (Kinga et al., 2015; Loshchilov & Hutter, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$. For the LLaMA-130M model, we use the same learning rate but reduce the mini-batch size to 32 with gradient accumulation to a global batch size of 512, train for 20,000 steps with 2,000 warm-up steps, and maintain the same optimizer configuration. For the LLaMA-350M model, we set the mini-batch size to 16 with gradient accumulation to a global batch size of 512, extend training to 60,000 steps with 6,000 warm-up steps. For the LLaMA-1B model, we set the mini-batch size to 16 with gradient accumulation to a global batch size of 512, extend training to 100,000 steps with 10,000 warm-up steps.
For the GPT-2 series, we follow the Sophia setup (Liu et al., 2023) with the nanoGPT implementation (Karpathy, 2022) and pretrain the vanilla GPT-2 series on OpenWebText (Gokaslan & Cohen, 2019). We set the maximum input sequence length to 1024 and adopt the bfloat16 precision. We use a global batch size of 480, cosine learning rate decay with 2,000 warm-up iterations, and global gradient clipping with a threshold of 1.0. All models are trained for 100,000 steps. We employ the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$ and weight decay of 0.1. For GPT-2 Small (124M), We use a per-device mini-batch size of 12 and 10 gradient accumulation steps across 4 GPUs, resulting in an effective global batch size of 480. We employ a cosine learning rate scheduler with a peak learning rate of 0.0006 and a minimum learning rate of 0.00003. For GPT-2 Medium (355M), we use a per-device mini-batch size of 10 and 12 gradient accumulation steps across 4 GPUs, resulting in an effective global batch size of 480. We employ a cosine learning rate scheduler with a peak learning rate of 0.0003 and a minimum learning rate of 0.00006.
For the hyperparameter noise factor in NormalNorm, we perform a lightweight Sobol-based search on the interval $[0,1]$, and set the value to $1$ for all experiments. For the hyperparameter $\lambda$ in each variant of IBNorm, we conduct a grid search over the set $\{2,3,4,5\}$. The final configurations are $\lambda=3$ for IBNorm-S, and $\lambda=4$ for both IBNorm-L and IBNorm-T. We investigated several hyperparameter configurations, including learning rate, learning rate scheduler, weight decay, and minibatch size, across all models and found the present configurations to generally work best.
@@ F.2 Vision Experiments
ResNet Experiments. In all experiments involving ResNet18 on CIFAR-10, we use stochastic gradient descent (SGD) with learning rate 0.1, weight decay 0.0005, momentum 0.9, and batch size 128 for LayerNorm and NormalNorm, with a noise factor of 0.4. For IBNorm experiments, we use learning rate 0.05 and weight decay 0.001. Following Eftekhari & Papyan (2025), models were trained from random initialization for 200 epochs, with a StepLR scheduler reducing the learning rate by a factor of 10 every 60 epochs.
For experiments with ResNet-50 on ImageNet, a batch size of 256 was used, with SGD parameters: learning rate 0.1, momentum 0.9, Nesterov acceleration enabled, and weight decay 0.0001, along with a cosine annealing learning rate scheduler with maximum 200 epochs and minimum learning rate 0.000001 for BatchNorm and NormalNorm, with a noise factor of 1. For IBNorm experiments, we use learning rate 0.01 and weight decay 0.001. For the hyperparameter $\lambda$ in each variant of IBNorm, we conduct a grid search over the set $\{2,3,4,5\}$. The final configurations are $\lambda=3$ for IBNorm-S, and $\lambda=4$ for both IBNorm-L and IBNorm-T. We explored several hyperparameter configurations, including learning rate, learning rate scheduler, weight decay, and minibatch size, across all models and found the present configurations to generally work best.
@@
We also follow the setup in NormalNorm (Eftekhari & Papyan, 2025) to train a ViT model with 8 transformer layers, 8 attention heads, hidden dimension size 768, MLP dimension size 2304, and patch size 16. Models were trained on ImageNet for 200 epochs with a global batch size of 512. Weighted random sampling based on inverse class frequency was applied during training to improve performance across all model configurations. The AdamW optimizer was used with learning rate 0.001, weight decay 0.05, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and epsilon 0.00000001 for LayerNorm, RMSNorm, and NormalNorm, with a noise factor of 1. For IBNorm, we set the learning rate as 0.0002. In addition, we employ a hybrid scheduling strategy that combines linear warmup with cosine annealing over 5 warmup training epochs. During the warmup phase, the learning rate increases linearly from 0.1 to 1.0 of the base learning rate. After warmup, a cosine annealing schedule is applied for the remaining epochs, with the learning rate decaying towards a minimum value of 0.000001.

### v2

@@ E.4 Proof for Lemma 5
$|\mathbb{E}[p(\tilde{T}_{1}|X)]-\mathbb{E}[p(\tilde{T}_{1}^{\prime}|X)]|\leq\delta,\quad|\mathbb{E}[p(\tilde{T}_{1}|Y)]-\mathbb{E}[p(\tilde{T}_{1}^{\prime}|Y)]|\leq\delta,$ (66)
$|\mathbb{V}(p(\tilde{T}_{1}|X))-\mathbb{V}(p(\tilde{T}_{1}^{\prime}|X))|\leq\delta,\quad|\mathbb{V}(p(\tilde{T}_{1}|Y))-\mathbb{V}(p(\tilde{T}_{1}^{\prime}|Y))|\leq\delta.$ (67)
@@ E.5 Proof for Corollary 2
With probability at least $1-\delta$ over training set $S$ of size $M<\infty$, the generalization error of a L-layer network $f_{\text{o}}$ satisfies
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U^{\text{o}}_{l}=\sqrt{\tfrac{-\text{IB}(T_{l}^{\text{o}})+C_{S}+\log(\frac{2}{\delta})}{M}},$
where $T^{\text{o}}_{l}$ is the intermediate representation at layer $l$, and $C_{S}$ is a term depending on the cardinality of the training set $S$. Here $f_{\text{o}}$ can be the network $f_{\text{S}}$ using standard normalization like LN and BN, or the network $f_{\text{IB}}$ with IBNorm.
@@
Because $\eta$ is linear and invertible, it does not affect the mutual information terms (Hjelm et al., 2018; Ash and Doléans-Dade, 2000); therefore, it suffices to analyze the intermediate representations before normalization representation recovery $\eta$: $\tilde{T}_{l}^{\text{o}}:=\psi_{l}^{o}\circ\zeta_{l}^{o}\circ h_{l}^{o}(X),\ l=1,\cdots,L$. Extending the results of Appendix F.4 in (Kawaguchi et al., 2023), we have: for arbitrary any $\delta>0$, with probability at least $1-\delta$ over the training set $S$ of size $M<\infty$, the following generalization bound holds:
$\text{gen}(S;f_{\text{o}})\;\leq\;\sum_{l=1}^{L}U_{l}^{\text{o}}:=\sum_{l=1}^{L}\sqrt{\tfrac{I(X;\tilde{T}_{l}^{\text{o}}|Y)+C(S,f_{l}^{\text{o}})+\log(\frac{2}{\delta})}{M}},$ (68)
where $\tilde{T}_{l}$ is the intermediate representation at layer $l$ before normalization representation recovery $\eta$ and $C(S,f_{l}^{\text{o}})\propto\ln(2|\Phi^{\text{o}}_{l}||\mathcal{Y}|/\delta)$ is a term depending on the cardinality of the hypothesis space associated with the l-th layer $f_{l}^{\text{o}}$ and the training set $S$. Since $f_{l}^{\text{IB}}$ and $f_{l}^{\text{S}}$ have the same parameter size, the model-complexity terms satisfy $C(S,f_{l}^{\text{IB}})\equiv C(S,f_{l}^{\text{S}})$ for all $l$. Thus, for the L-layer network, we denote $C_{S}:=C(S,f_{l}^{\text{IB}})\equiv C(S,f_{l}^{\text{S}})$. Thus, we obtain a bound on the total generalization error:
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U^{\text{o}}_{l}=\sqrt{\tfrac{I(X;\tilde{T}_{l}^{\text{o}}|Y)+C_{S}+\log(\frac{2}{\delta})}{M}}.$ (69)
Given the assumption that $\beta=1$, we have
$I(X;\tilde{T}_{l}^{\text{o}}|Y)\;=\;I(X;\tilde{T}_{l}^{\text{o}})-I(Y;\tilde{T}_{l}^{\text{o}})\;=\;-\text{IB}(\tilde{T}_{l}^{\text{o}}).$ (70)
Thus, we yield
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U^{\text{o}}_{l}=\sqrt{\tfrac{-\text{IB}(T_{l}^{\text{o}})+C_{S}+\log(\frac{2}{\delta})}{M}}.$ (71)
This establishes a layer-wise mutual information bound on the generalization error shown in Corollary 2. ∎
Remark. In our experiments, all models are trained on the same dataset $S$, so any dataset-dependent contribution to the generalization bound can be treated as invariant across different normalization schemes. The normalization layers considered in our study (LN, BN, and IBNorm) have identical total parameter size. Thus, the hypothesis spaces $\Phi^{\text{IB}}_{l}$ and $\Phi^{\text{S}}_{l}$ associated with $f^{\text{IB}}_{l}$ and $f^{\text{S}}_{l}$ have the same cardinality across all layers $l$. Theoretically, by Lemma 6 in (Kawaguchi et al., 2023), for the L-layer network, we have $C_{S}:=C(S,f_{l}^{\text{IB}})\equiv C(S,f_{l}^{\text{S}})$, that is, the model-complexity terms are equivalent for $f^{\text{IB}}_{l}$ and $f^{\text{S}}_{l}$. The generalization bound of $f_{\text{IB}}$ and $f_{\text{S}}$ therefore differ only in the representation-complexity term $\text{IB}(T_{l}^{\text{o}})\equiv I(X;\tilde{T}_{l}^{\text{o}}|Y)$ across all layers $l$.
@@ F.1 LLM Experiments
For all LLaMA models, we follow the setup in (Zhao et al., 2024) and pretrain the vanilla LLaMA series models from scratch on the C4 dataset. We set the maximum input sequence length to 512 and output sequence length to 256, and adopt the bfloat16 precision. For the LLaMA-60M model, we use a learning rate of $0.001$, mini-batch size of 64 with gradient accumulation to a global batch size of 512, 20,000 training steps with 1,000 warm-up steps, and the AdamW (Kinga et al., 2015; Loshchilov and Hutter, 2017) optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$. For the LLaMA-130M model, we use the same learning rate but reduce the mini-batch size to 32 with gradient accumulation to a global batch size of 512, train for 20,000 steps with 2,000 warm-up steps, and maintain the same optimizer configuration. For the LLaMA-350M model, we set the mini-batch size to 16 with gradient accumulation to a global batch size of 512, extend training to 60,000 steps with 6,000 warm-up steps. For the LLaMA-1B model, we set the mini-batch size to 16 with gradient accumulation to a global batch size of 512, extend training to 100,000 steps with 10,000 warm-up steps.
For the GPT-2 series, we follow the Sophia setup (Liu et al., 2023) with the nanoGPT implementation (Karpathy, 2022) and pretrain the vanilla GPT-2 series on OpenWebText (Gokaslan and Cohen, 2019). We set the maximum input sequence length to 1024 and adopt the bfloat16 precision. We use a global batch size of 480, cosine learning rate decay with 2,000 warm-up iterations, and global gradient clipping with a threshold of 1.0. All models are trained for 100,000 steps. We employ the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$ and weight decay of 0.1. For GPT-2 Small (124M), We use a per-device mini-batch size of 12 and 10 gradient accumulation steps across 4 GPUs, resulting in an effective global batch size of 480. We employ a cosine learning rate scheduler with a peak learning rate of 0.0006 and a minimum learning rate of 0.00003. For GPT-2 Medium (355M), we use a per-device mini-batch size of 10 and 12 gradient accumulation steps across 4 GPUs, resulting in an effective global batch size of 480. We employ a cosine learning rate scheduler with a peak learning rate of 0.0003 and a minimum learning rate of 0.00006.
For the hyperparameter noise factor in NormalNorm, we perform a lightweight Sobol-based search on the interval $[0,1]$, and set the value to $1$ for all experiments. For the hyperparameter $\lambda$ in each variant of IBNorm, we conduct a grid search over the set $\{2,3,4,5\}$. The final configurations are $\lambda=3$ for IBNorm-S, and $\lambda=4$ for both IBNorm-L and IBNorm-T. We investigated several hyperparameter configurations, including learning rate, learning rate scheduler, weight decay, and minibatch size, across all models and found the present configurations to generally work best. The statistical significance results are evaluated across three independent runs.
@@ F.2 Vision Experiments
ResNet Experiments. In all experiments involving ResNet18 on CIFAR-10, we use stochastic gradient descent (SGD) with learning rate 0.1, weight decay 0.0005, momentum 0.9, and batch size 128 for LayerNorm and NormalNorm, with a noise factor of 0.4. For IterNorm experiments, we use weight decay 0.005. For IBNorm experiments, we use weight decay 0.001. Following (Eftekhari and Papyan, 2025), models were trained from random initialization for 200 epochs, with a StepLR scheduler reducing the learning rate by a factor of 10 every 60 epochs.
For experiments with ResNet-50 on ImageNet, a batch size of 256 was used, with SGD parameters: learning rate 0.1, momentum 0.9, Nesterov acceleration enabled, and weight decay 0.0001, along with a cosine annealing learning rate scheduler with maximum 200 epochs and minimum learning rate 0.000001 for BatchNorm and NormalNorm, with a noise factor of 1. For IBNorm experiments, we use weight decay 0.001. For the hyperparameter $\lambda$ in each variant of IBNorm, we conduct a grid search over the set $\{2,3,4,5\}$. The final configurations are $\lambda=3$ for IBNorm-S, and $\lambda=4$ for both IBNorm-L and IBNorm-T. We explored several hyperparameter configurations, including learning rate, learning rate scheduler, weight decay, and minibatch size, across all models and found the present configurations to generally work best.
@@
We also follow the setup in NormalNorm (Eftekhari and Papyan, 2025) to train a ViT model with 8 transformer layers, 8 attention heads, hidden dimension size 768, MLP dimension size 2304, and patch size 16. Models were trained on ImageNet for 200 epochs with a global batch size of 512. Weighted random sampling based on inverse class frequency was applied during training to improve performance across all model configurations. The AdamW optimizer was used with learning rate 0.001, weight decay 0.05, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and epsilon 0.00000001 with a noise factor of 1. In addition, we employ a hybrid scheduling strategy that combines linear warmup with cosine annealing over 5 warmup training epochs. During the warmup phase, the learning rate increases linearly from 0.1 to 1.0 of the base learning rate. After warmup, a cosine annealing schedule is applied for the remaining epochs, with the learning rate decaying towards a minimum value of 0.000001.